Carlos and Robert leave Cabanatuan city at the same time. They are heading to Bulacan. Carlos, driving his sports car, travels 65kph. Robert, on a motorcycle travels at 55 kph. How long will it be before they are 55km apart? (show your solution)
Found 2 solutions by mananth, robertb:
Answer by mananth(16946)

Carlos and Robert leave Cabanatuan city at the same time. They are heading to Bulacan. Carlos, driving his sports car, travels 65kph. Robert, on a motorcycle travels at 55 kph. How long will it be before they are 55km apart? (show your solution)
...
Carlos, 65kph.
Robert, 55 kph.
How long will it be before they are 55km apart?
...
they are travelling in same dirction. every hour they distance themselves by 65-55 = 10 miles
rate = 10 mph
..
distance 55 km apart
t=d/r
t=55/10
t=5.5 hours.
They will be 55 km apart after 5.5 hours
...
m.ananth@hotmail.ca

Answer by robertb(5830)

Why are Carlos and Robert heading to Bulacan from Cabanatuan City?